Aleutian Company produces two products: Rings and Dings. They are manufactured in two departments: Fabrication and Assembly. Data for the products and departments are listed below.


Product
Number of
Units
Direct Labor Hours
Per Unit
Machine Hours
Per Unit
  Rings 1,010 3 7
  Dings 2,140 4 10

All of the machine hours take place in the Fabrication Department, which has estimated total factory overhead of $89,300. All of the labor hours take place in the Assembly Department, which has estimated total factory overhead of $65,900.

Aleutian Company uses the multiple production department factory overhead rate method. The Fabrication Department uses machine hours as an allocation base, and the Assembly Department uses direct labor hours.

The total factory overhead allocated per unit of Rings is

a.$49.27

b.$39.05

c.$54.16

d.$7.04
Answer #1

Total machine hour = (1010*7+2140*10) = 28470

Total labor hour = (1010*3+2140*4) = 11590

Fabrication overhead rate = 89300/28470 = 3.14 per machine hour

Assembly overhead rate = 65900/11590 = 5.69 per labor hour

Total factory overhead allocated per unit of Rings = (3*5.69+7*3.14) = 39.05

So answer is b) $39.05